OpenStudy (anonymous):

Partial fraction decomposition? 1/[x(x^2+1)] I got as far as 1= Ax^2 + A + Bx, and when I try to write the new system, I get A = 0 A = 1 B = 0 Which I know doesn't make sense, because A can't be both, right? What am I doing wrong?

hartnn (hartnn):

if there is a quadratic in denominator, we take cx+d/(ax^2+bx+c) so in your case , the partial fraction will be \(\large \dfrac{1}{x(x^2+1)}=\dfrac{A}{x}+\dfrac{Bx+c}{x^2+1}\) now try :)

hartnn (hartnn):

@MGR whenever you come online.

OpenStudy (amistre64):

one thing that always got me was when we had just an x^2 in a denominator, which is technically a quadratic, but it is treated as a linear factor - prolly because it is not prime if there is a prime quadratic, we use a linear on top.

OpenStudy (amistre64):

one neat thing to try is to factor it into its complex parts and just use constants for tops :)

hartnn (hartnn):

even for x^2 we use same thing, Bx+C, but we write it different way \(\large \dfrac{1}{x^2(x+1)}= \dfrac{A}{x+1}+\dfrac{Bx+C}{x^2}\) this same thing we write as \(\large \dfrac{A}{x+1} +\dfrac{B}{x}+\dfrac{C}{x^2}\)

OpenStudy (amistre64):

hmm, that does make sense ... i should prolly get some better textbooks :) thnx

OpenStudy (anonymous):

Thanks, hartnn! That was a lot better xD

hartnn (hartnn):

welcome ^_^